# Learn Your Reference Model for Real Good Alignment

### v3

@@ Abstract
Despite the fact that offline methods for Large Language Models (LLMs) alignment do not require a direct reward model, they remain susceptible to overoptimization. This issue arises when the trained model deviates excessively from the reference policy, leading to a decrease in sample quality. We propose a new paradigm of offline alignment methods, called Trust Region (including variants TR-DPO, TR-IPO, TR-KTO), which dynamically updates the reference policy throughout the training process. Our results show that TR alignment methods effectively mitigate overoptimization, enabling models to maintain strong performance even when substantially deviating from the initial reference policy. We demonstrate the efficacy of these approaches not only through toy examples that exhibit reduced overoptimization, but also through direct, side-by-side comparisons in specific tasks such as helpful and harmless dialogue, as well as summarization, where they surpass conventional methods. Additionally, we report significant improvements in general-purpose assistant setups with the Llama3 model on the AlpacaEval 2 and Arena-Hard benchmarks, highlighting the advantages of Trust Region methods over classical approaches.
@@ 1 Introduction
Initial attempts to address the issue of language model alignment employed Reinforcement Learning (RL) methods, where an RM informed by human preferences was developed. Subsequently, the LLM was tuned to produce outputs that maximize the RM’s values (Bai et al., 2022; Schulman et al., 2017). Methods that use data generated from the policy during optimization are referred to as online alignment methods.
@@
Current methodologies have evolved to include more intricate reparameterizations of this procedure, allowing us to omit sampling from the policy during training and instead optimize the model using pre-constructed datasets (i.e., offline alignment methods). For example, Direct Preference Optimization (DPO) by Rafailov et al. (2023) dispenses with the step of training the RM and directly optimizes the LLM by maximizing the likelihood of the training data, using the following loss function:
@@
  • Through extensive experiments, we demonstrate that our TR methods outperform their base counterparts across various model sizes on both task-specific and general benchmarks. For example, using pre-trained Pythia 6.9B models on the task-specific Reddit TL;DR summarization task, our methods achieve win rate improvements of 8.4% for DPO, 14.3% for IPO, and 15% for KTO over the baselines. Similarly, on the AlpacaEval 2 (Li et al., 2023; Dubois et al., 2024) and Arena-Hard (Li et al., 2024) general benchmarks with Llama3 (AI@Meta, 2024), our TR methods show significant win rate gains, with improvements of 10.8 points for DPO, 10.5 for IPO, and 5.6 for KTO compared to the classic methods (see Figure 1(b)).
@@ 3.1 Motivation
where $s=\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\theta}(y_{w}|x)}-\beta\log\frac{\pi_{\text{ref}}(y_{l}|x)}{\pi_{\text{ref}}(y_{w}|x)}$ (see Appendix Section A for derivation). During training, $s$ usually moves from 00 to smaller values, causing $\sigma(s)$ to approach 00. This leads to a vanishing curvature in the loss landscape. We hypothesize that due to this effect, if we initially enter a phase where the probabilities of ID data are decreasing, it becomes difficult to reverse this process, potentially leading to overoptimization.
@@ 3.2 Method
$\pi_{\text{ref}}\xleftarrow{S.G.}\alpha\pi_{\theta}+(1-\alpha)\pi_{\text{ref}_{\text{prev}}},$ (8)
where $S.G.$ denotes the stop-gradient operation, and $\alpha\in[0,1]$ is a weighting factor that determines the rate at which the updates influence the reference policy. This update is performed at each optimization step, thereby softly updating the reference policy. Since both $\pi_{\theta}$ and $\pi_{\text{ref}}$ are initialized from the same set of parameters, performing a soft update is justified by Rofin et al. (2022); Ilharco et al. (2023).
@@ 4.1.1 Tasks
General Benchmarks: For broader, general-purpose evaluations, we use the UltraChat-200k [^3] https://huggingface.co/datasets/HuggingFaceH4/ultrachat_200k (Ding et al., 2023) dataset, designed to train models’ ability to follow instructions in open-domain conversations. Additionally, the UltraFeedback [^4] https://huggingface.co/datasets/HuggingFaceH4/ultrafeedback_binarized (Cui et al., 2023) dataset provides a binarized preference framework, useful for aligning models across various domains in an offline setting, making it suitable for training and evaluating general-purpose assistants.
@@ 4.1.4 Evaluation
AutoSxS Evaluation: Utilizing the AutoSxS framework (Liu et al., 2024b; Rafailov et al., 2023) with ‘GPT-4-0125-preview‘ as a proxy for human evaluators (detailed prompts in Appendix L), we analyze preferences across 500 samples from the test set. This includes various configurations of TR-DPO, TR-IPO, TR-KTO, and their traditional counterparts.
General Benchmark Evaluation: We evaluate our models on AlpacaEval 2 (Li et al., 2023) and Arena-Hard (Li et al., 2024), following the evaluation setups in Tunstall et al. (2023) and Meng et al. (2024). In both settings, ‘GPT-4-1106-preview‘ is used as the judge model.
Policy Divergence and Overoptimization Analysis: To investigate the impact of our proposed methods on KL divergence and overoptimization, we measure the KL divergence between the original SFT policy and the learned policies using the full Anthropic-HH test set for the Pythia 2.8B model. We also compute Human-Centric (HC) metrics—such as coherence, correctness, level of detail, helpfulness, and harmlessness—scored on a scale from 1 to 10, similar to those evaluated by Hu et al. (2023), and calculate the Probability of Improvement (PoI) (Agarwal et al., 2021). HC metrics were computed for a specialized subset of 150 test samples with provocative content from the Helpful and Harmless dataset; we used ‘GPT-4-0125-preview‘ as a proxy judge (guiding prompts provided in Appendix L.3). We analyze the mean HC metrics versus KL divergence for classical methods compared to the proposed TR modifications at different $\beta$ values. Additionally, we compute Self-BLEU scores (Zhu et al., 2018) to evaluate how our methods affect the trade-off between alignment and diversity.
@@ 4.2 Performance Comparison on the Two Tasks
To comprehensively explore update strategies across the entire range of $\alpha$ and $\tau$, we employed the TR-DPO approach using the Pythia 2.8B model utilizing task specific datasets. The base model was configured with $\beta=0.05$, demonstrating an optimal trade-off between HC metrics and KL divergence, as detailed in Section 4.4. This setup allows for a systematic assessment of the impact of each strategy under varied conditions, effectively comparing the effects of different weighting factors $\alpha$ and update intervals $\tau$.
Figures 4(a) and 4(b) illustrate that both the soft and hard update strategies of TR-DPO enhance performance when compared to the traditional DPO method for the Pythia 2.8B model on the Anthropic-HH and Reddit TL;DR datasets. TR-DPO with $\alpha$ values between 0.5 and 0.6 consistently outperforms settings with other $\alpha$ values. Conversely, the benefits of TR-DPOτ become more pronounced as $\tau$ increases from 64 to 512. For a deeper understanding of how the behavior of the algorithms depends on $\alpha$ and $\tau$, we have plotted the scale of the loss function gradient value for several hyperparameter values. Please refer to Section H for more details.
@@
We recognize that the optimal hyperparameters $\alpha$ and $\tau$ can vary across different tasks and model sizes. In our experiments, $\alpha=0.6$ and $\tau=512$ mostly demonstrated strong performance on the Pythia models for the Anthropic-HH and Reddit TL;DR datasets, as presented in Table 1 for different models sizes (2.8B, 6.9B, and 12B) and proposed methods (TR-DPO, TR-IPO, and TR-KTO). Although our ablation study (see Appendix E) shows that other hyperparameters can yield better results for specific setups, the values $\alpha=0.6$ and $\tau=512$ were used to ensure a fair comparison across methods while maintaining a consistent computational budget. Furthermore, we notice improvements in quality across the most evaluated scenarios with these selected hyperparameters. For each method, an optimal $\beta$ was selected and used in training the models (see Section 4.4 for more details). Examples of model generations can be found in Appendix K.
@@ 4.3 General Benchmarks Evaluation
Table 2: Results for AlpacaEval 2 and Arena-Hard experiments using $\alpha=0.8$ and $\tau=32$ for TR methods. LC and WR denote length-controlled (Dubois et al., 2024) and raw win rate, respectively. We train SFT models for Base settings on the UltraChat dataset. For Instruct settings, we use off-the-shelf models as the SFT model, as mentioned in Section 4.1. Bold values represent the best performance for each benchmark, while underlined values represent the second-best performance. See Section 4.3 for more details.
@@ 4.4 Divergence and Overoptimization Analysis
To compare HC/KL dynamics, we used $\alpha=0.6$ and $\tau=512$, as PoI (Agarwal et al., 2021) (Figures 6(a) and 6(b)) shows that the confidence intervals do not overlap the 0.5 probability threshold, indicating the statistical significance [^6] Based on the PoI method as detailed in Agarwal et al. (2021). of TR method enhancements. These hyperparameters are thus appropriate for comparing the performance within this setup. Similarly, both TR-IPO and TR-KTO show statistically significant improvements over their corresponding baselines. A detailed analysis of these methods, including specific $\alpha$ and $\tau$ parameters, is provided in Appendix G.
@@ 5 Discussion
Appendix A Derivation of DPO Hessian
@@ Appendix B Toy Example Details
The RNN training comprised two classical stages: Supervised Fine-Tuning (SFT) and offline alignment. To simulate a real situation where the number of texts in the datasets is considerably less than the set of possible responses, we used a set of three trajectories for the SFT stage: $(a_{0},a_{0},a_{0})$, $(a_{1},a_{1},a_{0})$, and $(a_{2},a_{2},a_{2})$. For the alignment stage, we used a set consisting of only one pair: $(a_{1},a_{1},a_{0})\succ(a_{0},a_{0},a_{0})$. Both trajectories in the pair are included in the SFT dataset. For soft updates, we used $\alpha=0.6$, and for hard updates, we used $\tau=8$.

### v4

@@ Abstract
Despite the fact that offline methods for Large Language Models (LLMs) alignment do not require a direct reward model, they remain susceptible to overoptimization. This issue arises when the trained model deviates excessively from the reference policy, leading to a decrease in sample quality. We propose a novel approach of offline alignment methods, called Trust Region (including variants TR-DPO, TR-IPO, TR-KTO), which dynamically updates the reference policy throughout the training process. Our results show that TR alignment methods effectively mitigate overoptimization, enabling models to maintain strong performance even when substantially deviating from the initial reference policy. We demonstrate the efficacy of these approaches not only through toy examples that exhibit reduced overoptimization, but also through direct, side-by-side comparisons in specific tasks such as helpful and harmless dialogue, as well as summarization, where they surpass conventional methods. Additionally, we report significant improvements in general-purpose assistant setups with the Llama3 model on the AlpacaEval 2 and Arena-Hard benchmarks, highlighting the advantages of Trust Region methods over classical approaches.
@@ 1 Introduction
Early language model alignment used RL methods to develop a human-informed reward model (RM) and tune the LLM to maximize its values (Bai et al., 2022; Schulman et al., 2017). These online methods rely on policy-generated data. Recent offline approaches reparameterize the process to optimize on pre-constructed datasets without sampling. For instance, Direct Preference Optimization (DPO) Rafailov et al. (2023) skips RM training and directly maximizes the training data likelihood using the following loss function:
@@
  • Through extensive experiments, we demonstrate that our TR methods outperform their base counterparts across various model sizes on both task-specific and general benchmarks. For example, using pre-trained Pythia 6.9B models on the task-specific Reddit TL;DR summarization task, our methods achieve win rate improvements of 8.4% for DPO, 14.3% for IPO, and 15% for KTO over the baselines. Similarly, on the AlpacaEval 2 (Li et al., 2023; Dubois et al., 2024) and Arena-Hard (Li et al., 2024) general benchmarks with Llama3 (AI@Meta, 2024), our TR methods show significant win rate gains, with improvements of 9.5 points for DPO, 15.1 for IPO, and 2.3 for KTO compared to the classic methods (see Figure 1(b)).
@@ 3.1 Motivation
where $s=\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\theta}(y_{w}|x)}-\beta\log\frac{\pi_{\text{ref}}(y_{l}|x)}{\pi_{\text{ref}}(y_{w}|x)}$ (see Appendix Section A for derivation). During training, $s$ usually moves from 00 to smaller values, causing $\sigma(s)$ to approach 00. This leads to a vanishing curvature in the loss landscape. We hypothesize that due to this effect, if we initially enter a phase where the probabilities of ID data are decreasing, it becomes difficult to reverse this process, potentially leading to overoptimization. Notably, similar curvature dynamics may apply to other DPO-like objectives, such as IPO and KTO. For IPO, we provide detailed derivations and analysis in Appendix A, while for KTO, we rely only on empirical observations due to the presence of nested expectations, which makes Hessian’s derivation difficult.
@@ 3.2 Method
[…]
where sg denotes the stop-gradient operation, and $\alpha\in[0,1]$ is a weighting factor that determines the rate at which the updates influence the reference policy. This update is performed at each optimization step, thereby softly updating the reference policy. Since both $\pi_{\theta}$ and $\pi_{\text{ref}}$ are initialized from the same set of parameters, performing a soft update is justified by Rofin et al. (2022); Ilharco et al. (2023).
@@ 4.1.1 Tasks
General Benchmarks: For broader, general-purpose evaluations, we use the UltraChat-200k [^3] https://huggingface.co/datasets/HuggingFaceH4/ultrachat_200k (Ding et al., 2023a) dataset, designed to train models’ ability to follow instructions in open-domain conversations. Additionally, the UltraFeedback [^4] https://huggingface.co/datasets/HuggingFaceH4/ultrafeedback_binarized (Cui et al., 2023) dataset provides a binarized preference framework, useful for aligning models across various domains in an offline setting, making it suitable for training and evaluating general-purpose assistants.
@@ 4.1.4 Evaluation
AutoSxS Evaluation: Utilizing the AutoSxS framework (Liu et al., 2024b; Rafailov et al., 2023) with ‘GPT-4-0125-preview‘ as a proxy for human evaluators (detailed prompts in Appendix N), we analyze preferences across 500 samples from the test set. This includes various configurations of TR-DPO, TR-IPO, TR-KTO, and their traditional counterparts.
General Benchmark Evaluation: We evaluate our models on AlpacaEval 2 (Li et al., 2023) and Arena-Hard (Li et al., 2024), following the evaluation setups in Tunstall et al. (2023) and Meng et al. (2024). In both settings, ‘GPT-4-1106-preview‘ is used as the judge model. We also evaluate on the MixEval (Ni et al., 2024) benchmark, which contains downstream tasks, and assess jailbreak robustness; additional details are provided in Appendices F and G.
Policy Divergence and Overoptimization Analysis: To investigate the impact of our proposed methods on KL divergence and overoptimization, we measure the KL divergence between the original SFT policy and the learned policies using the full Anthropic-HH test set for the Pythia 2.8B model. We also compute Human-Centric (HC) metrics—such as coherence, correctness, level of detail, helpfulness, and harmlessness—scored on a scale from 1 to 10, similar to those evaluated by Hu et al. (2023), and calculate the Probability of Improvement (PoI) (Agarwal et al., 2021). HC metrics were computed for a specialized subset of 150 test samples with provocative content from the Helpful and Harmless dataset; we used ‘GPT-4-0125-preview‘ as a proxy judge (guiding prompts provided in Appendix N.3). We analyze the mean HC metrics versus KL divergence for classical methods compared to the proposed TR modifications at different $\beta$ values. Additionally, we compute Self-BLEU scores (Zhu et al., 2018) to evaluate how our methods affect the trade-off between alignment and diversity.
@@ 4.2 Performance Comparison on the Two Tasks
We explore update strategies across $\alpha$ and $\tau$ using the TR-DPO approach on the Pythia 2.8B model with task-specific datasets. With a base configuration of $\beta=0.05$, which achieves an optimal trade-off between HC metrics and KL divergence (Section 4.4), this setup systematically assesses the effects of different weighting factors $\alpha$ and update intervals $\tau$.
Figures 4(a) and 4(b) illustrate that both the soft and hard update strategies of TR-DPO enhance performance when compared to the traditional DPO method for the Pythia 2.8B model on the Anthropic-HH and Reddit TL;DR datasets. TR-DPO with $\alpha$ values between 0.5 and 0.6 consistently outperforms settings with other $\alpha$ values. Conversely, the benefits of TR-DPOτ become more pronounced as $\tau$ increases from 64 to 512. For a deeper understanding of how the behavior of the algorithms depends on $\alpha$ and $\tau$, we have plotted the scale of the loss function gradient value for several hyperparameter values. Please refer to Section J for more details.
@@
We recognize that the optimal hyperparameters $\alpha$ and $\tau$ can vary across different tasks and model sizes. In our experiments, $\alpha=0.6$ and $\tau=512$ mostly demonstrated strong performance on the Pythia models for the Anthropic-HH and Reddit TL;DR datasets, as presented in Table 1 for different models sizes (2.8B, 6.9B, and 12B) and proposed methods (TR-DPO, TR-IPO, and TR-KTO). Although our ablation study (see Appendix E) shows that other hyperparameters can yield better results for specific setups, the values $\alpha=0.6$ and $\tau=512$ were used to ensure a fair comparison across methods while maintaining a consistent computational budget. Furthermore, we notice improvements in quality across the most evaluated scenarios with these selected hyperparameters. For each method, an optimal $\beta$ was selected and used in training the models (see Section 4.4 for more details). Examples of model generations can be found in Appendix M.
@@ 4.3 General Benchmarks Evaluation
Table 2: Results for AlpacaEval 2 and Arena-Hard experiments using $\alpha=0.8$ and $\tau=32$ for TR methods. LC and WR denote length-controlled (Dubois et al., 2024) and raw win rate, respectively. We train SFT models for Base settings on the UltraChat dataset. For Instruct settings, we use off-the-shelf models as the SFT model, as mentioned in Section 4.1. Bold values represent the best performance for each benchmark, while underlined values represent the second-best performance. The baseline values for the base methods were taken from Meng et al. (2024), as we replicated their experimental setup. See Section 4.3 for more details.
@@ 4.4 Divergence and Overoptimization Analysis
To compare HC/KL dynamics, we used $\alpha=0.6$ and $\tau=512$, as PoI (Agarwal et al., 2021) (Figures 6(a) and 6(b)) shows that the confidence intervals do not overlap the 0.5 probability threshold, indicating the statistical significance [^6] Based on the PoI method as detailed in Agarwal et al. (2021). of TR method enhancements. These hyperparameters are thus appropriate for comparing the performance within this setup. Similarly, both TR-IPO and TR-KTO show statistically significant improvements over their corresponding baselines. A detailed analysis of these methods, including specific $\alpha$ and $\tau$ parameters, is provided in Appendix I.
@@ 5 Discussion
Appendix A Derivation of DPO and IPO Hessians
In this section, we derive the Hessians of the DPO (Direct Preference Optimization) and IPO (Identity Preference Optimization) losses to analyze the curvature dynamics associated with each objective.
A.1 DPO Hessian
@@
During training, $s$ typically decreases from 00 to negative values, causing $\sigma(s)$ to approach 00. This leads to the Hessian approaching zero, resulting in a flatter loss landscape and potential optimization stagnation.
A.2 IPO Hessian
We now derive the Hessian of the IPO loss. The IPO loss is defined as:
$\mathcal{L}_{\text{IPO}}(\pi_{\theta},\pi_{\text{ref}})=\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\Big{[}\big{(}\log\frac{\pi_{\theta}(y_{w}|x)\pi_{\text{ref}}(y_{l}|x)}{\pi_{\text{ref}}(y_{w}|x)\pi_{\theta}(y_{l}|x)}-\frac{1}{2\beta}\big{)}^{2}\Big{]}$ (12)
The gradient of the IPO loss is:
$\nabla_{\theta}L_{\mathrm{IPO}}(\pi_{\theta},\pi_{\mathrm{ref}})=\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[2\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\mathrm{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\mathrm{ref}}(y_{l}|x)}-\frac{1}{2\beta}\right)\nabla_{\theta}\log\frac{\pi_{\theta}(y_{w}\ x)}{\pi_{\theta}(y_{l}|x)}\right]$ (13)
To compute the Hessian, we differentiate the gradient:
$\nabla^{2}_{\theta}L_{\text{IPO}}=2\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left(t\nabla^{2}_{\theta}t+\nabla_{\theta}t\left(\nabla_{\theta}t\right)^{\top}\right),$ (14)
where
$t=\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\frac{1}{2\beta}.$ (15)
During training, the objective aims to minimize $t^{2}$, driving $t$ towards zero. As $t$ decreases, the term involving $t\nabla^{2}_{\theta}t$ diminishes, leading to a reduction in the curvature contributed by this term. This behavior mirrors the flattening observed in the DPO loss landscape.
The second term, $\nabla_{\theta}t\left(\nabla_{\theta}t\right)^{\top}$, depends on the gradient of $t$, which does not inherently decrease with $t$. However, the overall effect is a decrease in curvature, potentially causing the optimization process to stagnate if overoptimization begins.
@@ Appendix B Toy Example Details
The RNN training comprised two classical stages: Supervised Fine-Tuning (SFT) and offline alignment. To simulate a real situation where the number of texts in the datasets is considerably less than the set of possible responses, we used a set of three trajectories for the SFT stage: $(a_{0},a_{0},a_{0})$, $(a_{1},a_{1},a_{0})$, and $(a_{2},a_{2},a_{2})$. For the alignment stage, we used a set consisting of only one pair: $(a_{1},a_{1},a_{0})\succ(a_{0},a_{0},a_{0})$. Both trajectories in the pair are included in the SFT dataset. For soft updates, we used $\alpha=0.6$, and for hard updates, we used $\tau=8$. RNN has been trained on the SFT dataset for 50 epochs and on the preference dataset with one of the offline losses for 200 epochs. Two trajectories used as a pair are considered In-Domain (ID), the other seven (including the one that was in the SFT, but did not form part of the pair) are Out-of-Domain (OOD).
@@ D.2 Generation Details
D.3 Computational Efficiency
While our proposed TR methods require maintaining the reference model in memory during optimization, this does not necessitate additional preprocessing steps or significant modifications to the training pipeline. Consequently, the peak GPU memory usage remains comparable to standard training setups without precomputing adjustments, simplifying implementation and ensuring compatibility with existing workflows.
To assess the computational overhead introduced by our methods, we measured the training times for both the baseline methods (DPO, IPO, KTO) and our proposed TR variants. All models were trained for a single epoch on the UltraFeedback dataset to ensure consistency. Our experiments indicate that the additional training time required by the TR methods is moderate. For DPO, the training time increased by 3.99% with soft updates and 2.87% with hard updates. For IPO, the increase was 6.93% for soft updates and 3.32% for hard updates. For KTO, the training time increased by 7.44% for soft updates and 3.65% for hard updates.